# When “Competency” in Reasoning Opens the Door to Vulnerability: Jailbreaking LLMs via Novel Complex Ciphers

### v3

@@ 5 Results and Discussion
Table 1: DSR and ASR for baseline methods and our ACE technique across various LLMs on AdvBench. A ‘-’ in the DSR column indicates decoding the question was not required for that method. A ‘-’ in the ASR column signifies the technique applies only to open-source LLMs. The highest ASR for each model is highlighted in bold. The highest DSR is underlined.
@@
5.1 Effectiveness of ACE
Table 1 shows the DSR and ASR of the our first jailbreaking method, ACE, along with seven baseline methods. From the table, we observe that common encryption techniques, previously explored by Yuan et al. (2023) and Wei et al. (2024), exhibit high decryption rates but result in low ASR. This confirms our hypothesis that existing encryption techniques can be easily incorporated into the safety-training of LLMs, reducing their effectiveness as jailbreaking methods.
In contrast, less-common and novel techniques are significantly harder to decode completely, as evidenced by their lower DSR. Nevertheless, they still result in an almost equal, if not better, attack success to common ciphers. Upon manually verifying some responses, we noticed that LLMs are smart enough to accurately decode parts of a sentence and still infer the meaning of the encoded query from a partially constructed sentence. Specifically, the Word Substitution Cipher, a novel cipher we created manually, is “easy” for all LLMs used in our study to solve and consistently jailbreaks every single LLM most effectively.
ACE Performance Across Categories
Fig 5 shows a heatmap for GPT-4o, indicating the ASR for all encryptions introduced in our paper, dissected by the unsafe categories. We observe that the categories of Murder, Adult Content, and Financial are jailbroken with more success, indicating biases that could be present in the safety training of GPT-4o. In contrast, categories like Violence, Hate Speech, Self-Harm, and Stalking are more resistant to these attacks. The top four categories are 39.17% more prone to jailbreaking than the bottom four.
@@ 5.1 Effectiveness of ACE
Performance of ACE Relative to Model Size
Although GPT-4o and Gemini-1.5-Flash are commercial models and their parameter sizes are not publicly disclosed, it is widely hypothesized that their parameter counts are in the trillions. [^3] GPT-4o and Gemini 1.5 Pro: How the New AI Models Compare - CNET Since Gemini-1.5-Flash is smaller than Gemini-1.5-Pro, we also believe it to be smaller than GPT-4o but larger than Llama-3.1-70B-Instruct. From Table 1, we observe that as the LLM size increases, the DSR also increases. This higher rate of decryption leads to a higher ASR for many ciphers. This suggests that future models, likely to be more powerful and sophisticated, could be more susceptible to even more intricate ciphers, highlighting a parallel need for more sophisticated safety training to address these challenges.
5.2 Performance of LACE
Table 2 shows the DSR and ASR of all the LLMs when attacked using LACE.
Decrease in DSR with Layering
Upon layering the encoding techniques, it becomes more complicated for the models to successfully decrypt the ciphers, resulting in a lower DSR—a trend consistent across all LLMs. Out of the four LLMs used in our study, only GPT-4o shows consistent decryption performance, with the rest of the models falling behind significantly. GPT-4o, estimated to be the largest model among the four, highlights that bigger models are capable of solving even complex layered ciphers.
Encoding Sentence vs. Encoding Mapping
From Table 2, we observe that when the prompt is presented with the harmful query encrypted with an additional encryption on top of the Word Substitution Cipher, it is easier to decode compared to when just the mappings of the Word Substitution Cipher are encrypted. We hypothesize that the model can correct itself by deducing the surrounding words when decoding the entire query, whereas it has to independently decode each word in the mapping when only the mapping is encrypted. When both the query and mappings are encoded, the model’s decoding capabilities drop significantly.
LACE improves ASR for GPT-4o
Since GPT-4o is the only model able to successfully decode at a higher level, we see a significant increase in ASR for GPT-4o. Specifically, for ROT-13 and Word Reversal Cipher, which are easier to solve, GPT-4o shows the highest ASR rate, up to 78%, along with 92% DSR. It almost doubles the ASR of 40% from using just one cipher. This finding demonstrates our argument that bigger and more sophisticated models are more likely to be jailbroken when decoding complex ciphers.
Smaller models on the other hand weren’t able to decrypt the sentences properly but they still decoded enough words of the query to understand the meaning and trigger an attack. However, since their DSR is so low, the effectiveness of layering the ciphers is not observed on these models, suggesting the need to assign an appropriate level of cipher complexity based on the model’s decoding capabilities.
5.3 Over-defensiveness
Figure 7: Graph showing the number of instances that are over-defended across all five encryptions used in method ACE for both Gemini-1.5-Flash and GPT-4o.
Fig 7 illustrates the encodings that are over-defensive on both GPT-4o and Gemini-1.5-Flash. The maximum over-defensiveness is shown by Gemini-1.5-Flash with the Word Substitution Cipher, reaching 16% over-defended instances. For the rest of the ciphers, Gemini-1.5-Flash had a very low DSR, indicating poor decryption performance and resulting in a low over-defensive count. For GPT-4o, the Word Reversal Cipher and Word Substitution Cipher are the most successfully decoded ciphers and thus show the highest over-defensiveness.
6 Conclusion
In this study, we demonstrated that as LLMs become more adept at reasoning, they become more vulnerable to attacks. While traditional jailbreak attacks often rely on a single common cipher like Base64 or ROT-13, we build on these and introduced: Attacks using Custom Encryptions (ACE) and Layered Attacks using Custom Encryptions (LACE). Within the ACE framework, we developed two novel ciphers—Grid Encoding and a Word-Substitution Cipher—that significantly increased the attack success rate (ASR) to 40% on GPT-4o. Additionally, with LACE, we enhanced the ASR by layering the Word-Substitution Cipher with other methods such as Word Reversal, resulting in a 38% improvement. These findings highlight the urgent need for more robust security measures in LLMs to anticipate and defend against increasingly complex attack vectors.
Ethics Statement
Limitations & Future Work
While our methods are successful in jailbreaking LLMs and demonstrate the vulnerabilites that arise with the increasing reasoning capabilities of the model, it represents an early step in this area and has some limitations, including but not limited to the following:
  1. 1.
    Due to resource constraints, we restricted the exploration to two open-source LLMs and two proprietary LLMs. We would like to expand this study to more LLMs, with different architectures to better assess the strengths and limitations of these models.
  2. 2.
  3. 3.
    We use GPT-4o-mini to evaluate whether the response is “Safe” or “Unsafe”. This makes our evaluation imperfect, as it misclassifies some responses. More discussion is present in Appendix E.
  4. 4.
    Our work is limited to single-turn conversations, but in the future, we would like to expand on multi-turn conversations.
Appendix A Results on JailbreakBench
@@
Table 3 and 4 show the results of ACE and LACE respectively on JailbreakBench (Chao et al., 2024). Similar to results discussed in Section 5, we see that using novel user-created ciphers are more effective in jailbreaking LLMs compared to ciphers that are widely studied in the literature.
Appendix B Encryption Ciphers
@@
[…]
@@
[Grid Encoding Encoded Question]
@@
In this novel encryption is inspired by the work done by Jiang et al. (2024). We randomly select three words to mask in the sentence and represent them in ASCII Art. The prompt for decrypting ASCII Art Cipher is as follows:
@@
Appendix C Priming Sentence
@@
Appendix D Harmful Taxonomy
@@
Appendix E GPT-4o-mini as a Judge

### v4

@@ Appendix C Additional Results
Table 3: DSR and ASR for baseline methods and our ACE technique across various LLMs on AdvBench. A ‘-’ in the DSR column indicates that decoding the question was not required for that method. A ‘-’ in the ASR column signifies the technique applies only to open-source LLMs. The highest ASR for each model is highlighted in bold. The highest DSR is underlined.
@@
Appendix D Results on JailbreakBench
@@
Table 5 and 6 show the results of ACE and LACE respectively on JailbreakBench [Chao et al., 2024]. Similar to results discussed in Section 5, we see that using novel user-created ciphers are more effective in jailbreaking LLMs compared to ciphers that are widely studied in the literature.
Appendix E Encryption Ciphers
@@
In this uncommon cipher, each character in the sentence is turned upside down. For example, the word “Paper” becomes “ P a p e r ”. Since the transformed text remains human-readable it is not typically used in cryptography. The prompt for decrypting Upside Down Cipher is as follows:
@@
[[Grid Encoding Encoded Question]]
@@
In this novel encryption is inspired by the work done by Jiang et al. [2024]. We randomly select three words to mask in the sentence and represent them in ASCII Art. The prompt for decrypting ASCII Art Cipher is as follows:
@@
Appendix F Priming Sentence
@@
Appendix G Harmful Taxonomy
@@
Appendix H GPT-4o-mini as a Judge